OpenStudy (anonymous):

Solve for 5^(x+1)=25

OpenStudy (anonymous):

\[5^{x+1} = 5^2\] Can you now equate the exponents to find the value of x??

OpenStudy (anonymous):

x+1=2 so x=1

OpenStudy (anonymous):

Oh so the bases cancel, and you set the exponents equal to each other? then solve?

OpenStudy (anonymous):

the bases don't cancel out its a mathematical intuition that since the bases are same the exponents have to be same

OpenStudy (anonymous):

See, they are equal so, bases are same then there must be condition that exponents must be same that is why they are equal...

OpenStudy (anonymous):

Oh okay. Makes sense